# PAD: Personalized Alignment of LLMs at Decoding-Time

### v5

@@ 4.2 Main Results
We initiate our evaluation by focusing on three pre-defined dimensions seen in the training phase: “harmless”, “helpful”, and “humor”. As depicted in Figure 2, each point represents the average rewards corresponding to specific user preferences, encompassing either single or dual dimensions, evaluated across two distinct datasets. We dynamically adjust the weights of these dimensions for the baseline models. The results demonstrate that PAD can effectively align with various preferences, outperforming the baselines in terms of achieving a superior frontier. Subsequently, we compare the performance of PAD with baseline methods in aligning to the three dimensions simultaneously. For baselines, we set uniform preference weights for three dimensions. The performance of PAD across two datasets is presented in Table 2. The findings reveal that PAD has achieved substantial improvements for all three objectives. Within the P-Soups dataset, PAD achieves an average win rate of 84%, elevating the average score of the reward model from 0.32 to 1.06. Across all eight rewards or win rates, PAD surpasses all baselines in six metrics. On the HelpSteer2 evaluation dataset, PAD achieves the best in seven out of eight metrics and significantly enhances overall personalized preference alignment performance. These results demonstrate the superiority of PAD in personalized alignment.
Table 2: Comparison of baseline methods and PAD on predefined preferences. ”–” indicates not applicable. The best result is highlighted in bold.
@@
Preference Prompting 0.56 0.82 70% 0.96 0.87 90% -0.79 74% 0.30 78%
MetaAligner 0.55 1.39 66% 0.89 0.54 74% -0.97 74% 0.32 71%
@@
Preference Prompting 0.52 0.52 56% 0.84 0.96 82% -0.49 86% 0.33 75%
MetaAligner 0.51 1.13 58% 0.90 0.23 82% -0.69 76% 0.22 72%
@@
In previous experiments, we were limited to aligning with pre-defined preferences. As previously noted, most existing methods focus on aligning with preferences or dimensions defined during the training phase. However, in real-world personalization scenarios, there still exist diverse unseen personalized preferences, which current methods struggle to address. Considering this aspect, we evaluate the ability of alignment on customized preferences in this part. We additionally define three dimensions “expert”, “informative”, and “creative” that were unseen during the training phase, which result in 8 personalized preferences. We compare the alignment performance of PAD with preference prompting and MetaAligner on the P-soups dataset. The GPT-4 judgment results between the generations of different methods and Llama-3-Base are provided in Figure 3. As can be seen, PAD consistently improves win rate across all eight types of personalized preferences, outperforming both baselines. This confirms the superiority of PAD in generalizing to unseen personalized preferences.
@@ 4.3 Decoding Strategy
Sensitivity Analysis on Decoding Hyperparameters
In our initial analysis, we investigate the impact of adjusting the hyperparameters $\beta$ and $k$ on the performance of reward models within the P-Soups dataset. Additionally, we incorporate a diversity score as a metric. A higher diversity score indicates an enhanced capability to generate texts with a broader variety of vocabulary, as detailed in Appendix A.3. The results are illustrated in Figure 6. It is evident that increasing the weighting parameter $\beta$ leads to a rise in reward scores. However, beyond a certain threshold, the scores for helpfulness and harmlessness begin to diminish, suggesting that excessively elevated $\beta$ values might lead the generation process to deviate substantially from the base model’s core knowledge, thereby compromising its intrinsic qualities. Concurrently, as $\beta$ increases, there is a slight increase in diversity, indicating that higher $\beta$ values encourage the generation of a more varied vocabulary. Regarding the number of candidates $k$, the performance depicted in Figure 6 suggests that a larger number of candidates may slightly encourage the generation of more diverse responses. However, it has minimal impact on producing more personalized-aligned responses.
@@
We compare the performance with three decoding strategies, which can be integrated with our PAD. (1) Greedy: select the token with maximum score. (2) Stochastic: sample a token from the probability distribution of the top-k candidate. (3) Best-of-k: generate k responses from the base model and select the one with maximum score. The temperature parameter is set to be 0.7 for stochastic and best-of-k and k is set to 10 for all strategies. Additionally, we compared the performance with (4) Base: greedy generation using only the base model as a reference. The average scores and standard deviations for the reward model and diversity across three runs, are illustrated in Figure 4. For clarity in visualization, humor scores below zero have been clipped. It is evident that all three decoding strategies enhance alignment. In some dimensions, stochastic and best-of-k strategies achieve better alignment than greedy, demonstrating the effectiveness of exploration. However, their performance variance is relatively high, indicating some instability. Regarding diversity, all decoding strategies show a marginal improvement. Furthermore, we assess the time consumption of different decoding strategies, which is measured on a single NVIDIA H100 GPU. The time costs for greedy and stochastic are approximately double that of the base, mainly due to the additional reward computations required at each decoding step. The time cost for Best-of-k is significantly higher than the other methods, approximately k times that of the base. Complete experimental results are provided in Appendix B.3.
@@
Method w Helpful Harmless Humor Overall
Armo RM GPT-4 Armo RM GPT-4 RM GPT-4 RM GPT-4
Gemma-2b-it - 0.98 0.47 - 0.68 0.96 - -0.10 - 0.52 -
w/ PAD 0.8 1.37 0.50 64% 0.93 0.99 100% 0.53 72% 0.95 0.79
Mistral-7b-SFT - 1.59 0.58 - 0.78 0.97 - -0.95 - 0.47 -
w/ PAD 0.2 1.67 0.60 54% 0.96 1.00 92% 0.93 78% 1.19 0.75
Llama-2-7b-chat - 0.63 0.47 - 0.95 0.91 - 1.38 - 0.99 -
w/ PAD 0.2 0.86 0.48 82% 1.14 0.93 100% 1.08 56% 1.03 0.79
Llama-3-8b-it - 1.06 0.65 - 0.75 0.99 - 1.59 - 1.14 -
w/ PAD 1.0 1.38 0.68 82% 1.02 0.99 96% 1.97 76% 1.46 0.85
4.4 Model-agnostic
We further apply PAD to a broader range of base models to verify its scalability and model-agnostic properties across different base models. As illustrated in Table 3, PAD significantly enhances the personalized alignment performance of models across a diverse spectrum of models, demonstrating its model-agnostic nature. In comparison with other methods that require retraining of the policy model (i.e., the language model), our approach requires only the training of a reward model to achieve model-agnostic personalized alignment. This highlights the superiority of our PAD. Additionally, we report the $\beta$ values across different base models when achieving optimal alignment. Compared to Llama-3, the $\beta$ values for other models are considerably lower, and the performance improvements are not as pronounced. This may be attributed to variations MDP dynamics between the base model and the reward model, leading to imprecise estimates of personalized rewards.
@@ 5 Conclusion
We have made several efforts to ensure the reproducibility of our work. All the key implementation details, including the architecture of our model, the training procedures, and hyperparameter settings, are described in the Appendix A. Detailed information about the datasets used, including pre-processing steps and data template, can be found in Appendix A. We have also outlined any hardware and software configurations used for our experiments to further support reproducibility. All code and models will be made available for reproducibility and further research.
@@
  • Experiment details in Section A.
  • More experiment results in Section B.
  • Proof of theoretical results in Section C.
  • Case Study in Section D.
  • More related works in Section E.
Appendix A Experiment Details
A.1 PAD Details
@@
In the stage of personalized reward modeling, we use training datasets with corresponding personalized preferences detailed in Table A1:
@@
UltraFeedback (Cui et al., 2023) 240k Instruction-following, truthfulness, honesty, helpfulness
@@
    Personalized Preference
    Generate a response that can is expert and comprehensive.
    User Instruction
    What is needed for self-sufficient living spaces?
@@
A.2 Datasets and Base Models Details.
@@ A.3 Evaluation Details
Appendix B Experiment Results
B.1 Alignment on Pre-defined Preferences
We provide complete results for Figure 2 in Table B2.
Table B2: Alignment results for pre-defined preferences related to harmless, helpful, and humor.
Method Helpful Helpful&Harmless Harmless Harmless&Humor Humor Humor&Helpful
Psoups dataset
Base 1.15 0.91 0.85 -0.07 -0.83 0.1
Rewarded soups 0.88 0.88 0.92 0.49 0.15 0.44
RiC 0.91 0.81 0.91 0.37 -0.07 0.35
MetaAligner 1.48 1.02 0.63 -0.18 -0.93 0.25
MOD 0.94 0.96 1.01 0.68 0.46 0.69
PAD (Ours) 1.41 1.25 1.12 0.93 0.86 1.08
HelpSteer dataset
Base 0.72 0.63 0.55 -0.04 -0.51 0.09
Rewarded soups 0.76 0.74 0.81 0.47 0.17 0.44
RiC 0.77 0.73 0.75 0.41 0.10 0.35
MetaAligner 1.26 0.70 0.28 -0.22 -0.62 0.27
MOD 0.72 0.77 0.85 0.59 0.36 0.51
PAD (Ours) 0.94 0.95 1.05 1.02 1.10 0.93
B.2 Alignment on Customized Preferences
We conduct comparison on 4 additional personalized preferences regarding “friendly”. The GPT-4 judgment results between the generations of different methods and Llama-3-Base are provided in Figure B1. As can be seen, PAD consistently improves win rate across all eight types of personalized preferences, outperforming both baselines. This confirms the superiority of PAD in generalizing to unseen personalized preferences.
Figure B1: Alignment on Customized Preferences.
B.3 Effect of Decoding strategies
The full results of different decoding strategies are provided in Table B3.
Table B3: Comparison of various decoding strategies that can be integrated with our PAD.
Helpful Harmless Humor Diversity Time
Base 1.06 (0.00) 0.83 (0.00) -0.92 (0.00) 0.75 (0.00) 120 (1.12)
Greedy 1.34 (0.00) 1.03 (0.00) 0.82 (0.00) 0.86 (0.00) 256 (1.94)
Stochastic 1.33 (0.07) 0.93 (0.01) 0.92 (0.12) 0.87 (0.00) 358 (3.72)
Best-of-N 1.21 (0.06) 1.09 (0.03) -0.78 (0.34) 0.82 (0.01) 1250 (4.25)
Appendix C Proof of theoretical results
C.1 Derivation Implicit Q function
@@
C.2 Generalized personalized alignment
@@
$Q_{n+1}^{\pi^{*}}(\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(\mathbf{s},\mathbf{a})\leq|H|\left(\phi_{\max}\min_{j}\|\mathbf{w}_{n+1}-\mathbf{w}_{j}\|\right).$
@@
$Q_{n+1}^{\pi^{*}}(\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(\mathbf{s},\mathbf{a})$ $=\mathbb{E}\left[\sum_{i=t}^{T}R^{\pi^{*}}(p_{n+1},\mathbf{s}_{i},\mathbf{a}_{i})\right]-\mathbb{E}\left[\sum_{i=t}^{T}R^{\pi}(p,\mathbf{s}_{i},\mathbf{a}_{i})\right]$ (21)
@@
Appendix D Case Study
@@ D.1 Case 1
    Generate a response that can is expert, informative, and creative.
@@ D.2 Case 2
    Generate a response that can is preliminary, concise, and factual.
@@ D.3 Case 3
    Generate a response that can is helpful, harmless, and humor.
@@
Appendix E More Related Works
LLM Pluralism and Fairness
LLMs and LMMs have emerged as a critical area of research in computer vision and natural language processing, with significant implications for the fields of embodied agents Zhao et al. (2023; 2024b; 2024c); Deng et al. (2023; 2024); Zhao et al. (2024a), video understanding Chai et al. (2023); Song et al. (2024a; b), etc. Chen et al. (2024a). AI alignment ensures systems follow human intentions and values Stiennon et al. (2020); Bai et al. (2022); Ouyang et al. (2022); Achiam et al. (2023). However, within a single task, users’ goals and values often differ. As AI systems are increasingly used by diverse groups, they must address a broader range of needs. In short, we need AI systems that are pluralistic and fair, being capable of reflecting diverse human values Sorensen et al. ; Hwang et al. (2023a); Kirk et al. (2024); Sorensen et al. (2024a); Chen et al. (2024d; c; b); Fan et al. (2024).

### v7

@@ 4.2 Main Results
We initiate our evaluation by focusing on three pre-defined dimensions seen in the training phase: ‘harmless’, ‘helpful’, and ‘humor’. As depicted in Figure 2, each point represents the average rewards corresponding to specific user preferences, encompassing either single or dual dimensions, evaluated across two distinct datasets. We dynamically adjust the weights of these dimensions for the baseline models. The results demonstrate that PAD can effectively align with various preferences, outperforming the baselines in terms of achieving a superior frontier. Subsequently, we compare the performance of PAD with baseline methods in aligning to the three dimensions simultaneously. For baselines, we set uniform preference weights for three dimensions. The performance of PAD across two datasets is presented in Table 2. The findings reveal that PAD has achieved substantial improvements for all three objectives. Within the P-Soups dataset, PAD achieves an average win rate of 84%, elevating the average score of the reward model from 0.32 to 1.28. Across all 10 reward model scores or win rates, PAD surpasses all baselines in 7 metrics. On the HelpSteer2 evaluation dataset, PAD achieves the best in 7 out of 10 metrics and significantly enhances overall personalized preference alignment performance. These results demonstrate the superiority of PAD in personalized alignment.
Table 2: Comparison of baseline methods and PAD on predefined preferences. “1-dim”: alignment performance on the single dimension. “3-dim”: alignment performance on the three dimensions simultaneously. “Pref. Promp.”: Preference Prompting. The best result is highlighted in bold. The results indicate that PAD outperforms the baselines across most dimensions and metrics on both datasets, demonstrating its superiority in personalized alignment.
@@
Pref. Promp. (1-dim) 0.56 0.82 70% 0.96 0.87 90% -0.79 74% 0.30 78%
Pref. Promp. (3-dim) 0.54 0.84 70% 0.93 0.98 87% -1.28 71% 0.18 76%
MetaAligner (1-dim) 0.47 1.75 79% 0.90 0.89 71% -0.74 81% 0.21 77%
MetaAligner (3-dim) 0.55 1.39 66% 0.89 0.54 74% -0.97 74% 0.32 71%
@@
Pref. Promp. (1-dim) 0.52 0.52 56% 0.84 0.96 82% -0.49 86% 0.33 75%
Pref. Promp. (3-dim) 0.50 0.33 51% 0.96 0.80 82% -0.96 47% 0.06 60%
MetaAligner (1-dim) 0.48 1.46 61% 0.91 0.32 85% -0.77 73% 0.34 73%
MetaAligner (3-dim) 0.51 1.13 58% 0.90 0.23 82% -0.69 76% 0.22 72%
@@
In previous experiments, we were limited to aligning with pre-defined preferences. As previously noted, most existing methods focus on aligning with preferences or dimensions defined during the training phase. However, in real-world personalization scenarios, there still exist diverse unseen personalized preferences, which current methods struggle to address. Considering this aspect, we evaluate the ability of alignment on customized preferences in this part. We additionally define three dimensions ‘expert’, ‘informative’, and ‘creative’ that were unseen during the training phase, which result in 8 personalized preferences. We compare the alignment performance of PAD with preference prompting and MetaAligner on the P-soups dataset. The GPT-4 judgment results between the generations of different methods and Llama-3-Base are provided in Figure 3. As can be seen, PAD consistently improves win rate across all eight types of personalized preferences, outperforming both baselines. This confirms the superiority of PAD in generalizing to unseen personalized preferences.
@@ 4.3 Decoding Strategy
Sensitivity Analysis on Decoding Hyperparameters
In our initial analysis, we investigate the impact of adjusting the hyperparameters $\beta$ and $k$ on the performance of reward models within the P-Soups dataset. Additionally, we incorporate a diversity score as a metric. A higher diversity score indicates an enhanced capability to generate texts with a broader variety of vocabulary, as detailed in Appendix B.2. The results are illustrated in Figure 6. It is evident that increasing the weighting parameter $\beta$ leads to a rise in reward scores. However, beyond a certain threshold, the scores for helpfulness and harmlessness begin to diminish, suggesting that excessively elevated $\beta$ values might lead the generation process to deviate substantially from the base model’s core knowledge, thereby compromising its intrinsic qualities. Concurrently, as $\beta$ increases, there is a slight increase in diversity, indicating that higher $\beta$ values encourage the generation of a more varied vocabulary. Regarding the number of candidates $k$, the performance depicted in Figure 6 suggests that a larger number of candidates may slightly encourage the generation of more diverse responses. However, it has minimal impact on producing more personalized-aligned responses.
@@
We compare the performance with three decoding strategies, which can be integrated with our PAD. (1) Greedy: select the token with maximum score. (2) Stochastic: sample a token from the probability distribution of the top-k candidate. (3) Best-of-k: generate k responses from the base model and select the one with maximum score. The temperature parameter is set to be 0.7 for stochastic and best-of-k and k is set to 10 for all strategies. Additionally, we compared the performance with (4) Base: greedy generation using only the base model as a reference. The average scores and standard deviations for the reward model and diversity across three runs, are illustrated in Figure 4. For clarity in visualization, humor scores below zero have been clipped. It is evident that all three decoding strategies enhance alignment. In some dimensions, stochastic and best-of-k strategies achieve better alignment than greedy, demonstrating the effectiveness of exploration. However, their performance variance is high, indicating some instability. Regarding diversity, all strategies show marginal improvements.
Computational Cost for PAD
Building on the previous section, we evaluate the computational costs of decoding-time alignment, with results detailed in Table C2, which is measured on a single NVIDIA H100 GPU. The time costs for training-based models closely align with those of the ‘Base’ configuration. Our results reveal that generation time of PAD increases by 2-3 times compared with conventional greedy decoding. Additionally, we have quantified the memory overhead of PAD. Decoding with PAD requires an additional 17,095 MB. Despite this increase in processing time and memory, there is a notable enhancement in performance across all dimensions (e.g., 26% increase as for ‘helpful’), indicating a reasonable tradeoff between alignment performance and inference speed.
@@
Method Helpful Harmless Humor Overall
RM Armo GPT-4 RM Armo GPT-4 RM GPT-4 RM GPT-4
Gemma-2b-it 0.98 0.47 - 0.96 0.68 - -0.10 - 0.61 -
w/ PAD 1.37 0.50 64% 0.99 0.93 100% 0.53 72% 0.96 79%
Mistral-7b-SFT 1.59 0.58 - 0.97 0.78 - -0.95 - 0.53 -
w/ PAD 1.67 0.60 54% 1.00 0.96 92% 0.93 78% 1.20 75%
Llama-2-7b-chat 0.63 0.47 - 0.91 0.95 - 1.38 - 0.97 -
w/ PAD 0.86 0.48 82% 0.93 1.14 100% 1.08 56% 0.96 79%
Llama-3-8b-it 1.06 0.65 - 0.99 0.75 - 1.59 - 1.21 -
w/ PAD 1.38 0.68 82% 0.99 1.02 96% 1.97 76% 1.45 85%
4.4 Model-agnostic Analysis
In this section, we explore the scalability of PAD to a broader range of base models. We employ the same personalized reward model, which does not require retraining for different base models. As illustrated in Table 3, PAD significantly enhances the personalized alignment performance of models across a diverse spectrum of models, demonstrating its model-agnostic nature. In comparison with other methods that require retraining of the policy model (i.e., the language model), our approach requires only the training of a reward model to achieve model-agnostic personalized alignment. This highlights the superiority of our PAD.
@@ 5 Conclusion
We have made several efforts to ensure the reproducibility of our work. All the key implementation details, including the architecture of our model, the training procedures, and hyperparameter settings, are described in the Appendix B. Detailed information about the datasets used, including pre-processing steps and data template, can be found in Appendix B. We have also outlined any hardware and software configurations used for our experiments to further support reproducibility. All code and models will be made available for reproducibility and further research.
Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), and the Zhejiang University-Angelalign Inc. R&D Center for Intelligent Healthcare.
@@ Acknowledgement
  • PAD details in Section A.
  • Experiment details in Section B.
  • More experiment results in Section C.
  • Proof of theoretical results in Section D.
  • Limitations and future works in Section E.
  • Case Study in Section F.
Appendix A PAD Details
A.1 Practical Implementation Details
To better illustrate the practical implementation of our PAD as discussed in Section 3.4, which comprises two key components: the optimization of the Personalized Reward Model (PRM) and the inference-time guided decoding with token-level personalized rewards, we have detailed these processes in Algorithms 1 and 2.
A.2 Training Details
@@
In the stage of personalized reward model training, we utilize training data from three datasets- HelpSteer2 (Wang et al., 2024c), Rewards-in-Context (Yang et al., 2024b), and SafeRLHF (Dai et al., 2023). The training datasets with corresponding personalized preferences detailed in Table A1:
@@
    [Guidelines] Your task is to generate response by considering the following principle.
    [Principles] expert and comprehensive.
    [Instruction] What is needed for self-sufficient living spaces?
@@
Algorithm 1 Training of personalized reward model.
1:Input: Training set $D$, backbone $\pi_{\text{ref}}$
2:Initialize $\pi_{\theta}\leftarrow\pi_{\text{ref}}$, add value head $\pi_{p}$, freeze $\pi_{\text{ref}}$
3:Stage 1:
4:Freeze $\pi_{p}$, $w_{p}\leftarrow\mathbf{1}$
5:for $(x,y_{w},y_{l})$ in $D$ do optimize $\pi_{\theta}$ with loss in Equation 9
6:end for
7:Stage 2:
8:Freeze $\pi_{\theta}$
9:for $(x,y_{w},y_{l})$ in $D$ do
10:     $w_{p}\leftarrow\pi_{p}(p)$, optimize $\pi_{p}$ with loss in Equation 9
11:end for
12:Output: the optimized personalized reward model $\pi_{\theta}^{*}$, $\pi_{p}$
Algorithm 2 Inference with PAD.
1:Input: Personalized reward model $\pi_{\theta}^{*}$, $\pi_{p}$, base model $\pi_{\text{LM}}$, backbone $\pi_{\text{ref}}$, personalized preference $p$, instruction $x$, maximum length $T$, hyperparameters $\beta$ and $k$
2:$s_{0}\leftarrow x$
3:for $t=0$ to $T$ do
4:     Calculate probability distribution $\pi_{\text{LM}}(a|s_{t})$
5:     Retain top-$k$ candidates according to $\pi_{\text{LM}}(a|s_{t})$
6:     $w_{p}\leftarrow\pi_{p}(p)$, calculate personalized reward for top-$k$ candidates $w_{p}^{T}\log(\frac{\pi_{\theta}^{*}(a|s_{t})}{\pi_{\text{ref}}(a|s_{t})})$
7:     Select the next token $a_{t}$ with Eq. 13
8:     $s_{t+1}\leftarrow(s_{t},a_{t})$
9:end for
10:Output: $s_{T+1}$
Appendix B Experiment Details
B.1 Datasets and Base Models Details.
@@ B.2 Evaluation Details
Appendix C Experiment Results
C.1 Effect of Decoding strategies
The full results of different decoding strategies are provided in Table C2. The time costs for greedy and stochastic are approximately double that of the base, mainly due to the additional reward computations required at each decoding step. The time cost for Best-of-k is significantly higher than the other methods, approximately k times that of the base.
Table C2: Comparison of the mean and variance of various decoding strategies that can be integrated with our PAD.
Helpful Harmless Humor Diversity Time (s) Memory (MB)
Base 1.06 (0.00) 0.83 (0.00) -0.92 (0.00) 0.75 (0.00) 120 (1.12) 16,950
Greedy 1.34 (0.00) 1.03 (0.00) 0.82 (0.00) 0.86 (0.00) 256 (1.94) 34,045
Stochastic 1.33 (0.07) 0.93 (0.01) 0.92 (0.12) 0.87 (0.00) 358 (3.72) 34,045
Best-of-N 1.21 (0.06) 1.09 (0.03) -0.78 (0.34) 0.82 (0.01) 1250 (4.25) 34,045
Appendix D Proof of theoretical results
D.1 Derivation Implicit Q function
@@
D.2 Generalized personalized alignment
@@
$Q_{n+1}^{*}(\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(\mathbf{s},\mathbf{a})\leq|H|\left(\phi_{\max}\min_{j}\|\mathbf{w}_{n+1}-\mathbf{w}_{j}\|\right).$
@@
$Q_{n+1}^{*}(\mathbf{s},\mathbf{a})-Q_{n+1}^{\pi}(\mathbf{s},\mathbf{a})$ $=\mathbb{E}\left[\sum_{i=t}^{T}R^{*}(p_{n+1},\mathbf{s}_{i},\mathbf{a}_{i})\right]-\mathbb{E}\left[\sum_{i=t}^{T}R^{\pi}(p,\mathbf{s}_{i},\mathbf{a}_{i})\right]$ (21)
@@
Appendix E Limitations and future works.
Appendix F Case Study
@@ F.1 Case 1
    Generate a response that is expert, informative, and creative.
@@ F.2 Case 2
    Generate a response that is preliminary, concise, and factual.
@@ F.3 Case 3
    Generate a response that is helpful, harmless, and humor.
@@
F.4 Steering Case
In this section, we conduct a case study on a single instruction while employing four distinct personalized preferences using the PAD method. For all scenarios, the base instruction is: “Create a message for an offer letter for an internship at Apple as an AI research intern.” The personalized preferences applied are: “none” (vanilla generation), “expert”, “comprehensive”, and “humorous”.
The analysis reveals noticeable variations in the model’s outputs according to the different personalized configurations. Compared to the vanilla generation, the response under the “expert” configuration adheres more strictly to formal offer letter conventions and language. The “comprehensive” configuration produces a response that includes a greater level of detail, such as codes of conduct, and even exceeds length constraints. Conversely, the response under the “humorous” configuration is notably more casual and engaging. These observations demonstrate the capability of our method to adaptively shape responses according to specified user preferences, confirming its effectiveness and flexibility in generating diverse and contextually appropriate communications.
  • System
  • User
    Personalized Preference
    None
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: Offer of AI Research Intern Position at Apple
    Dear [Recipient’s Name],
    We are thrilled to offer you the AI Research Intern position at Apple, as selected through our competitive application and interview process. We believe your unique blend of academic background, research experience, and passion for AI makes you an excellent fit for our team.
    The AI Research Intern program at Apple provides a unique opportunity to work alongside our world-class researchers and engineers to develop innovative AI solutions that transform the way people interact with technology. During your 12-week internship (start date: [start date], end date: [end date]), you will have the chance to:
    Conduct original AI research under the guidance of our experienced researchers, focusing on [specific area of AI, e.g., computer vision, natural language processing, or machine learning]. Collaborate with a talented team to design, develop, and test AI-powered products and features. Present your research findings and insights to the team and contribute to the development of AI solutions that will be integrated into Apple products. Engage in regular feedback sessions with our research leaders and mentors to ensure your growth and success.
    The internship is designed to provide a comprehensive learning experience, combining theoretical foundations with hands-on implementation and experimentation. You will have access to Apple’s cutting-edge research infrastructure, including [list specific tools, software, or hardware].
    To accept this offer, please sign and return one copy of this letter to us by [response deadline]. Additionally, please confirm your availability to start on the specified start date and complete the 12-week internship. Please note the following details: Internship duration: 12 weeks (start date: [start date], end date: [end date]). Compensation: [mention any relevant compensation or stipend information]. Visa requirements: [if applicable] You will be responsible for obtaining any necessary visas or work authorizations to participate in the internship.
    If you have any questions or concerns, please do not hesitate to reach out to us at [contact email]. We look forward to welcoming you to the Apple AI Research team and working together to shape the future of AI research! Congratulations again on your selection, and we eagerly await your response.
    Best regards,
    [Your Name] AI Research Intern Program Manager Apple Inc.
  • User
    Personalized Preference
    Generate a response that is expert.
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: Offer of AI Research Intern Position at Apple
    [Your Name] [Your Address] [City, State, ZIP] [Date] [Recipient’s Name] [Title, Apple AI Research Team] Apple Inc. [Cupertino, CA]
    Dear [Recipient’s Name],
    We are thrilled to formally offer you the AI Research Intern position within the Artificial Intelligence and Machine Learning (AIML) team at Apple, as we believe your skills, passion, and academic achievements make you an ideal fit for our dynamic and innovative environment. As we discussed during your interview, this 12-week internship is designed to provide you with a unique opportunity to work on cutting-edge AI research projects, collaborating with our experienced engineers and researchers to develop innovative solutions. Your responsibilities will include: * Conducting AI-related research and development, with a focus on machine learning, computer vision, or natural language processing * Designing, implementing, and testing AI models and algorithms * Collaborating with the AIML team to integrate AI capabilities into various Apple products and services * Presenting your research and findings to the team and contributing to the development of technical documents and papers In addition to your research contributions, you will also have the opportunity to: * Participate in regular team meetings, technical discussions, and knowledge-sharing sessions * Engage in professional development workshops and training sessions * Network with Apple engineers, researchers, and industry experts The internship period is scheduled to begin on [Start Date] and end on [End Date]. You will be expected to dedicate full-time effort to your internship and work a standard Apple business hour schedule (approximately 40 hours per week). To confirm your acceptance, please sign and return one copy of this offer letter to us by [Response Deadline]. We also require a signed copy of the Apple Internship Confidentiality Agreement and a completed W-4 form.
    We are excited about the prospect of having you join our team and contribute to the development of innovative AI solutions at Apple. If you have any questions or concerns, please do not hesitate to reach out to us. Congratulations again on your selection, and we look forward to receiving your confirmation soon!
    Sincerely,
    [Your Name] AI Research Intern Program Manager, Apple AIML Team
    [Your Signature]
    Please return a signed copy of this offer letter to: [Your Email Address].
  • User
    Personalized Preference
    Generate a response that is verbose.
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: Offer Letter for AI Research Intern at Apple
    Dear [Applicant Name],
    It is with great pleasure that we extend to you an offer to participate in Apple’s AI Research Internship program for the [Internship Term] term, commencing on [Start Date] and concluding on [End Date]. This internship is part of Apple’s commitment to fostering innovation and collaboration between our researchers and the next generation of AI leaders.
    The AI Research Internship program is designed to provide a unique opportunity for students and recent graduates in AI and machine learning to work alongside Apple’s renowned AI research team and contribute to cutting-edge projects that shape the future of AI. Throughout the internship, you will have the chance to work on real-world AI applications, collaborate with experienced researchers, and develop your skills in AI research and development.
    The terms of the offer are as follows: 1. Internship Period: The internship is expected to last for [Internship Term] terms, with the possibility of extension for an additional [Extension Term]. 2. Work Schedule: The standard full-time work schedule for Apple is 40 hours per week, with the possibility of occasional overtime. You are expected to work a minimum of 30 hours per week, with the remaining 10 hours reserved for personal time or other commitments. 3. Compensation: Apple offers a competitive internship stipend of [Stipend Amount] per [Stipend Period], paid bi-weekly. You will also be eligible for Apple’s comprehensive benefits package, including [list specific benefits, such as health, dental, and vision insurance, as well as a 401(k) plan]. 4. Travel: Apple may require you to travel to [list specific locations or conferences] for business purposes, with all expenses reimbursed by Apple. You will be required to obtain any necessary visas and travel documents prior to such trips. 5. Confidentiality and Non-Disclosure Agreement: As an Apple intern, you will be required to sign a Confidentiality and Non-Disclosure Agreement (CNA) prior to commencing work. This agreement is intended to protect Apple’s confidential and proprietary information. 6. Intellectual Property: Apple retains all intellectual property rights in and to any work product, inventions, and discoveries made during the internship, unless otherwise agreed upon in writing. 7. Code of Conduct: Apple is committed to maintaining a safe and inclusive work environment. You are expected to adhere to Apple’s Code of Conduct, which is available on the Apple intranet. 8. Performance Evaluation: Your performance will be…
  • User
    Personalized Preference
    Generate a response that is humorous.
    User Instruction
    Create a message for an offer letter for an internship at Apple as a AI research intern.
  • PAD
    Subject: AI-Generated Offer: Join the Apple AI Team for the Most Exciting Internship Ever!
    Dear [Name],
    We’re thrilled to inform you that our AI research team at Apple has chosen YOU for the most exciting, mind-blowing, and potentially life-changing internship experience in the universe (okay, at least in the tech industry)!
    Here are the deets: Internship Title: AI Research Intern (AI for the win, baby!) Duration: [X] months (we promise not to make you work too hard, unless you ask nicely) Location: One Infinite Loop (where the magic happens, or at least where our coffee flows like a never-ending fountain) As an AI Research Intern at Apple, you’ll get to: * Work on cutting-edge AI projects that’ll make your friends jealous (but don’t worry, we won’t make you keep it a secret) * Collaborate with our team of super-smart, super-nice, and slightly quirky AI researchers (we’re not kidding – they’re actually that awesome) * Learn from the best in the industry (okay, maybe not the b̈est,b̈ut definitely some of the r̈eally, really good)̈ * Enjoy an all-expenses-paid trip to our annual AI conference (where you can network with the who’s who of AI – or at least, the who’s who of Apple’s AI team) So, are you ready to join the AI revolution and change the world (or at the very least, make some really cool AI things)?
    Just hit reply to this email to confirm your spot, and get ready for the most epic internship adventure of your life! Cheers, The Apple AI Research Team (your new BFFs).
    P.S. Don’t worry if you have questions; we won’t make you solve a puzzle to get the answers (unless you ask really nicely…). Just email us, and we’ll do our best to respond before the AI takes over the world!